OpenStudy (anonymous):

Can somebody tell me if I'm right? Given f(x) = 4x+1/3, solve for f-1(3) I got 2 The answer choices are 2 4 6 8

OpenStudy (freckles):

4x+1/3 or (4x+1)/3 ?

OpenStudy (anonymous):

4x+1 over 3 Kinda like a fraction :)

OpenStudy (anonymous):

Can you explain to me how it's 4? :)

OpenStudy (freckles):

so you mean (4x+1)/3 ? \[f^{-1}(3)=a \\ \text{ then } f(a)=3 \\ \text{ so solve } \frac{4a+1}{3}=3 \]

OpenStudy (freckles):

can you finish solving 4a+1=9

OpenStudy (anonymous):

Yes but there are no parenthesis

OpenStudy (anonymous):

Here

OpenStudy (freckles):

(4x+1)/3 means \[\frac{4x+1}{3}\] where as 4x+1/3 means \[4x+\frac{1}{3}\]

OpenStudy (freckles):

yes I made that assumption above

OpenStudy (freckles):

\[f^{-1}(3)=a \\ \text{ then } f(a)=3 \\ \text{ so solve } \frac{4a+1}{3}=3 \] I was asking you to solve this for a to verify your answer

OpenStudy (freckles):

multiply 3 on both sides 4a+1=9 and 2 more steps and you will have your answer

OpenStudy (anonymous):

4a=8 Divide by 4 =2?

OpenStudy (freckles):

yep